OpenStudy (anonymous):

find the L.C.M of 8,12 and 14

OpenStudy (anonymous):

2 * 4 = 8 2 * 6 = 12 2 * 7 = 14 can you find it now?

OpenStudy (anonymous):

is it two @gaara438125

OpenStudy (anonymous):

yeah :]